What is the standard form when your x intercept is -2 and your y intercept is 6

Answer by Alan3354(69443)

What is the standard form when your x intercept is -2 and your y intercept is 6
----------
What is the standard form when your x intercept is A and your y intercept is B
--> Bx + Ay = A*B
---------------
What is the standard form when your x intercept is -2 and your y intercept is 6
6x - 2y = -12
---
3x - y = -6